OpenStudy (anonymous):

sqrt(3)tanx+1=0 find all solutions

OpenStudy (raden):

sqrt(3)tanx+1=0 sqrt(3)tanx = -1 tanx = -1/sqrt(3) = -1/3 sqrt(3) solve for x (remember these is the special angles)

OpenStudy (anonymous):

thank you! I am just confused on how to solve for x on problems like these

OpenStudy (raden):

tanx = -1/3 sqrt(3) in the first quadran will satisfied for x = 150 degrees or 5pi/6 in the 4th quadrant will satisfied for x = 330 degrees or 11pi/6

OpenStudy (raden):

btw, what's interval for x given there ?

OpenStudy (anonymous):

i think its (0, 2pi)

OpenStudy (raden):

kay, that's all solution have done above

OpenStudy (anonymous):

thanks, so for the problem 2cos^2theta +cos theta=1 i factored it and got cos theta=1/2 and cos theata=-1. but from here how do u find the solutions for theta?

OpenStudy (raden):

the key is "Memorize the trigono values for the special angles" :D

OpenStudy (anonymous):

right. but when/how do i know when to subtract/add by either pi or 2 pi